Assume that you are a vending machine dealer. You plan to purchase a vending machine for $200,000. One year later, you are expected to sell it back and receive $224,000 as cash. What is the IRR (internal rate of return) on this investment? Be sure to apply the definition of IRR and show your work.

In order to buy the vending machine, you plan on borrowing $80,000 from Bank at 5% of interest rate and raise $120,000 from investors at 10% of cost of equity, are you going to accept this project? Why? (Tax rate is 30%). Show your work.
Answer #1

Hello,
In excel by applying formula
=IRR(values,[guess])

Year Amount
0 -200000
1 224000
IRR 12%


Screen shoot of excel



WACC - Wd*Kd + Wd*Ke
Kd = 5(1-tax)'
Kd = 5(1-0.3)
Kd = 3.5%

Ke = 10%
Weights = (cost of debt)Cost of borrowing = 80000 = 80000/200000 = 40%
Weights of equity = 1- .40
Weight of equity =60%

WACC = .4*3.5 + .6*10
WACC =7.4%

Yes, Project should be accepted beacuse IRR> WACC

I hope this clear your doubt.

Feel free to comment if you still have any query or need something else. I'll help asap.

Do give a thumbs up if you find this helpful.